OpenStudy (anonymous):

Two real-valued functions, f(x) and g(x), are defined as f(x) = x2 + 6x + 9 and g(x) = x2 – 9. Find the results of combining these functions. Tiles (f + g)(x) (f – g)(x) (f × g)(x) (f ÷ g)(x) Pairs 6(x + 3) 2x(x + 3) (x+3)(x−3) x4 + 6x3 – 54x – 81

OpenStudy (misty1212):

tiles?

OpenStudy (misty1212):

\[f(+g)(x)=f(x)+g(x)=x^2+x+9+x^2-9\] combine like terms others are similar to that one

OpenStudy (anonymous):

i have to match those tiles to the pairs. ex. (x+3)(x-3)=(f/g)(x)